Find the variance for the sample given in a.

2, 4, 7, 12, 15



1
Expert's answer
2022-03-02T10:08:12-0500

Sample variance can be found as

"s^2={\frac {\displaystyle\sum_{i=1}^n(x_i-x)^2} {n-1}}" , where n - sample size, x - sample mean

In the given case

"x={\frac {2+4+7+12+15} 5}=8"

"s^2={\frac {(2-8)^2+...+(15-8)^2} {5-1}}={\frac {118} 4}=29.5"